OpenStudy (korosh23):

Find the DOMAIN of the function in INTERVAL NOTATION. F(x) = cube root(5t-7) I will post my own attempt in a sec.

OpenStudy (korosh23):

\[f(x) = \sqrt[3]{5t-7}\] ---> This is the question My own attempt: \[5t-7 \ge 0\] \[t \ge 7/5\] Domain in interval notation: \[[7/5, +\infty )\]

OpenStudy (zzr0ck3r):

You only need to do this for even roots, cube root has all real numbers \(\mathbb{R}\) as domain.

OpenStudy (zzr0ck3r):

But if that 3 were a 4, you nailed it :)

OpenStudy (korosh23):

oh you are right . So it would be (-infinity, infinity).

OpenStudy (zzr0ck3r):

Correct

OpenStudy (korosh23):

Thanks a lot @zzr0ck3r :)

OpenStudy (zzr0ck3r):

This is one of those times that open study worked exactly how I would love it to always work. One little thing led you to doing the thing you would do for the even case, so you come here and write it up, obviously not just looking for answers, and in writing it up, I almost guarantee that you will not forget it come test day. awesome